Grade 12th pass10 grade science

Refractive index of a glass with respect to water is (9/8). Refractive index of a glass with respect to air is (3/2). Find the refractive index of water with respect to air.

Harshit Singh
4 Years agoGrade 12th pass

1 Answer

Yash Patari
4 Years ago
Dear student

Refractive index is defined as the ratio of the speed of light in a vacuum to its speed in a specific medium.

Let us consider

Refractive index of air = Ra

Refractive index of glass = Rg

Refractive index of water = Rw

The refractive index of glass with respect to water is given as,

Rgw = Rg/Rw

9/8 = Rg/Rw……………………………..(1)

The refractive index of glass with respect to air is

Rga = Rg/Ra……………………………..(2)

The refractive index of water with respect to air is

Rwa = Rw/Ra……………………………..(3)

Substitute the value of (1), (2) in (3) we get

Rwa = 4/3

The refractive index of water with respect to air is 4/3.